# Feasibility Study on Active Learning of Smart Surrogates for Scientific Simulations

### v1

@@ 4.2. Evaluation Metrics
Table 1. Comparison of test MAE between the prediction of the DNN surrogates and ground-truth simulations using U-Net (left) and CNN autoencoder (right) network for different percentages of labelled data across different acquisition functions. The metrics are measured for different lattice regions of interest as depicted in Figure 4. Red indicate measures better or similar to the random acquisition.
Figure 7. Comparison of absolute error maps between DNN predictions and the ground truth using between random (top row), diversity (middle row), and TOD (bottom row) acquisition function at different percentage of labelled data on U-Net architecture for two example lattice (1) and (2)
Figure 8. (a) Initial condition, (b) steady-state solution (ground truth), prediction of (c) U-Net and (e) CNN autoencoder surrogate, and prediction error of (d) U-Net and (f) CNN autoencoder surrogate, trained on 100% data, for two examples of two source simulation data where sources of 5-pixel radius are placed randomly within a 100 x 100 lattice. The scale for the error color bar for UNET and CNN autoencoder are different.
5. Experiments
We split the two-source data described in section 3.1. into 16,000 training data, 4,000 validation data, and 4,000 test data.
We trained the U-Net and CNN autoencoder structure defined in section 3.2. We divided the 16000 training data into an initial 1000 labelled data (samples with simulation outputs available) and a 15000 unlabelled pool (samples without simulation outputs). We performed active learning with the acquisition functions defined in section 4.1 for CNN autoencoder and exclude entropy acquisition function for U-Net due to large computational demand for ${}^{\prime}N^{\prime}$ forward passes. At each acquisition round, new unlabled samples are selected to query for simulation outputs, and the DNN surrogates were retrained with the current labeled dataset (input-output simulationp pairs). This was repeated until all the 15000 unlabelled samples are labeled. In each active learning round, the DNN surrogate was trained for 500 epochs and saved based on the best validation loss. For the entropy acquisition function, in CNN autoencoder, 40% dropouts were added after batch normalization of the first and second convolutional layers. The code repository can be found in (Bajracharya, 2023).
5.1. Comparison of Active Learning Acquisition Functions
Figure  5(a) compares weighted-MAE between the ground truth and DNN surrogate predictions for different acquisition functions at different percentages of labeled data, evaluated against random acquisitions for the U-Net architecture. As shown, actively selecting simulation data for DNN surrogates training, especially by considering the estimated loss of the surrogate (TOD), consistently improves the DNN performance compared to random acquisition using the same amount of labeled training data. Note that TOD, even though reliant on an estimate of the DNN loss without the simulation output as labels, acquires a performance on par with using the actual DNN loss calculated from labels as shown in Figure  6(a). In comparison, diversity-based selection of training samples demonstrated limited benefits with a performance similar to or marginally better than random acquisition.
Table  1 lists the weighted MAE between the simulated ground truth and predictions on the test data for U-Net across quarterly percentages of labeled data, computed for the different regions of interest as described in Figure  4. The results show that, at the same amount of labeled training data throughout the active learning process, the careful selection of training simulations achieved a test performance improvement across all acquisition strategies (except when all data are used): this improvement was more evident when the size of the labeled data was small. Additionally, these improvements were more pronounced at the regions proximal to the source (SRC, RING1) where the errors were the highest. This observation can be better appreciated in visual examples of the absolute errors (Figure  7) between DNN surrogate predictions and ground-truth simulation outputs obtained on the U-Net architecture. As expected, across the three rows, there is a gradual decrease in error as the training data increases. Notably, however, TOD consistently exhibits lower errors compared to random training, especially in regions close to the sources.
@@ 5.1. Comparison of Active Learning Acquisition Functions
Figure 9. Comparison of absolute error maps between DNN predictions and the ground truth using between random (top row), diversity (middle row), and TOD (bottom row) acquisition function at different percentage of labelled data on CNN autoencoder. The scale for the error color bar is different from U-Net in Figure  7
5.2. Effect of Architecture on Active Learning
@@
Figure  5(b) compares the weighted MAE for the CNN autoencoder at different percentages of labeled data across different acquisition functions. Compared with Figure  5(a), it is apparent that despite using identical acquisition functions and datasets, the performance improvements due to active learning is more pronounced in U-Net architecture compared to CNN autoencoder: the performance gain due to active learning for the CNN autoencoder is marginal at best with random training. This distinction is also evident in the quantitative data presented in Table  1. Similarly, as shown in the the visual examples of the absolute error maps between ground-truth simulations and surrogate predictions presented in Figure  9, in contrast to our previous observations on the U-Net, the errors among the acquisition functions remain consistently similar.
These results suggest an interesting and important finding: for active selection of training simulations to play a role in the construction of DNN surrogates, it is important to first identify an appropriate if not optimal DNN architecture for the surrogate, as the choices of architectures have a significant effect on the relative ranking of various acquisition function and the benefit they can deliver. Interestingly, this result is consistent with those reported systematic evaluations of general DAL methods (Beck et al., 2021; Munjal et al., 2022).
6. Discussion
@@
7. Conclusion
We present an investigative study that underscores the benefit of utilizing active learning in training a diffusion solver surrogate model. We experimentally show that for certain acquisition functions, active learning with fewer data (¡50% data) shows promise in improving over training randomly with larger data size (¿75% data). This provides a strong foundation for the next steps to build up the HPC infrastructure of Smart Surrogates where training simulations are generated on the fly as steered by active learning, potentially on a DNN architecture optimized for the scientific simulations at hand.
Acknowledgements.

### v2

@@ 5.1 Comparison of Active Learning Acquisition Functions
Table 1: Comparison of test MAE between the prediction of the DNN surrogates and ground-truth simulations using U-Net (left) and CNN autoencoder (right) network for different percentages of labelled data across different acquisition functions. The metrics are measured for different lattice regions of interest as depicted in Figure 4. Red indicate measures better or similar to the random acquisition.
@@
5.2 Effect of Architecture on Active Learning
@@
Figure 6(b) compares the weighted MAE for the CNN autoencoder at different percentages of labeled data across different acquisition functions. Compared with Figure 6(a), it is apparent that despite using identical acquisition functions and datasets, the performance improvements due to active learning is more pronounced in U-Net architecture compared to CNN autoencoder: the performance gain due to active learning for the CNN autoencoder is marginal at best with random training. This distinction is also evident in the quantitative data presented in Table  1. Similarly, as shown in the the visual examples of the absolute error maps between ground-truth simulations and surrogate predictions presented in Figure 9, in contrast to our previous observations on the U-Net, the errors among the acquisition functions remain consistently similar.
These results suggest an interesting and important finding: for active selection of training simulations to play a role in the construction of DNN surrogates, it is important to first identify an appropriate if not optimal DNN architecture for the surrogate, as the choices of architectures have a significant effect on the relative ranking of various acquisition function and the benefit they can deliver. Interestingly, this result is consistent with those reported systematic evaluations of general DAL methods [57, 58].
6 Discussion
@@
7 Conclusion
We present an investigative study that underscores the benefit of utilizing active learning in training a diffusion solver surrogate model. We experimentally show that for certain acquisition functions, active learning with fewer data (<50% data) shows promise in improving over training randomly with larger data size (>75% data). This provides a strong foundation for the next steps to build up the HPC infrastructure of Smart Surrogates where training simulations are generated on the fly as steered by active learning, potentially on a DNN architecture optimized for the scientific simulations at hand.
8 Acknowledgements